OpenStudy (anonymous):

If f(x) =x+2/2 , then what is f–1(x) ? x + 1 2x + 1 2x – 2 2x + 2

hero (hero):

y = (x+2)/2 x = (y+2)/2 2x = y +2 2x-2 = y f^-1(x) = 2x-2

hero (hero):

Third option, sorry